OpenStudy (anonymous):

The integral of Sin2x/Sinxdx

OpenStudy (anonymous):

Sin(2x) is a trig identity which expands out to 2sinxcosx.

OpenStudy (anonymous):

Okay then should I set u=sin2x? and dv=dx

OpenStudy (anonymous):

No. if sin2x = 2sinxcosx, you can cancel out the sinx

OpenStudy (anonymous):

Okay so I set u=2sinxcosx?

OpenStudy (anonymous):

There is no u substitution. sin2x is a trig identity sin2x is EQUAL to 2sinxcosx. This means that your equation can be rewritten as 2sinxcosx/sinx. You can cancel out the sinx So you end up with 2cosx And then take the integral of 2cosxdx

OpenStudy (anonymous):

Okay....Then u should be cosxdx?

OpenStudy (anonymous):

Where is u coming from?

OpenStudy (anonymous):

From the last part integral of 2cosxdx

OpenStudy (anonymous):

When I take the integral of that should u=cosx?

OpenStudy (anonymous):

Where is u? The integral is just \[\int\limits_{}^{}2cosxdx\]

OpenStudy (anonymous):

I need to find the anti-derivative to the integral of 2cosxdx

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So shouldnt I set u=cosx

OpenStudy (anonymous):

\[1/2\int\limits_,2cosxdx\] u=cosx, du=sinxdx

OpenStudy (anonymous):

You don't need u at all. \[\int\limits_{}^{}2cosxdx = 2\int\limits_{}^{}cosxdx\] Which equals 2sinx + c.